# Understanding LLMs’ Cross-Lingual Context Retrieval: How Good It Is And Where It Comes From

### v1

@@ A.1 Full list of models evaluated
A.2 Detailed evaluation results
@@
A.3 Prompt used in error type detection
The v1 prompt format is:
The v2 prompt format is:
The prompt for error type detection is:
<|im_start|>
user
@@ A.3 Prompt used in error type detection
A.4 Detailed language error and generation failure error rates
Tables 8 and 9 show detailed language and generation failure error rates across all tested languages. Meanwhile, Table 10 provides a more granular view of the generation failure errors discussed in the main text. While Table 9 presents the aggregated rate of these errors, Table 10 is further subdivided into three separate tables: Table 10(a), Table 10(b), and Table 10(c). These tables individually display the error rates for gibberish errors, refusal errors, and blank errors, respectively, across all tested models and languages in the 2-shot en-x xMRC task setting.
@@ A.4 Detailed language error and generation failure error rates
Table 9: 2-shot generation failure error rates (%) on en-x xMRC tasks.
@@
Table 10: Detailed percentages (%) of generation failure error types (gibberish error, refusal error, and blank error) on 2-shot en-x tasks.
Appendix B Two-phased xMRC Analysis
B.1 Further analysis on MRD
B.1.1 Example of Attribution
@@
B.1.2 MRD for other LLaMA models
Figures 12, 13 and 14 provide further illustrative examples of the mean MRD for context and question components, specifically for LLaMA-3.1-8B, LLaMA-3.1-Instruct-70B and LLaMA-2-Chat-7B. These figures complement the MRD analysis presented in the main body of this paper.
@@ B.1.2 MRD for other LLaMA models
B.1.3 Analysis of task descriptions and demonstrations
Analyzing the MRD of task descriptions and demonstrations in our 2-shot setting (Figures 15-17) reveals a general trend where demonstrations tend to exhibit a comparable or slightly higher MRD than task descriptions across the LLaMA model family, suggesting demonstrations are at least as important as, if not slightly more impactful than, task descriptions in guiding the models. This could indicate that providing concrete examples is a particularly effective way to communicate the desired behavior for cross-lingual context retrieval to these models.
@@ B.1.3 Analysis of task descriptions and demonstrations
B.1.4 Influence of prompt format on MRD pattern
@@
Figure 18: Mean MRD for LLaMA-3.1-Instruct-8B on both "balanced" and "en-superior" samples in v1 prompting format. Only the results of context and question parts of the prompt are displayed.
B.2 Hidden state similarity results for other LLaMA models
@@
B.3 Training details and evaluation results of our finetuned LLaMA-3.1-8B

### v2

@@ B.1 Full list of models evaluated
B.2 Detailed evaluation results
@@
B.3 Performance vs. parameter size
To observe the effect of increasing model size on xMRC performance, we evaluate the Qwen-2.5 models from 1.5B to 72B, showing results in Table 5. The trend is consistent with our findings in the main body, and an advantage of the 7B model stands out.
Version Size en-en mean en-x mean x-x en-x / en-en x-x / en-en
Base 1.5B 61.26 47.10 54.55 0.77 0.89
3B 71.09 58.75 60.00 0.83 0.84
7B 62.42 57.51 66.11 0.92 1.06
14B 76.71 69.04 74.60 0.90 0.97
32B 80.37 73.23 74.02 0.91 0.92
72B 86.03 78.92 81.16 0.92 0.94
Instruct 1.5B 70.63 55.21 58.11 0.78 0.82
3B 71.65 64.41 60.39 0.90 0.84
7B 81.83 76.43 71.61 0.93 0.88
14B 80.91 67.64 72.95 0.84 0.90
32B 80.97 73.77 71.90 0.91 0.89
72B 77.12 66.04 70.29 0.86 0.91
Table 5: xMRC performances of Qwen-2.5 models with increasing parameter sizes.
@@
B.4 Detailed language error and generation error rates
Tables 9 and 10 show detailed language and generation error rates across all tested languages. Note that Table 10 contain the results judged by both Qwen-2.5-Instruct-72B and Gemini-2.5-Flash-Lite, which are consistent to each other. Meanwhile, Table 11 provides a more granular view of the generation errors discussed in the main text.
While Table 10 presents the aggregated rate of these errors, Table 11 is further subdivided into three separate tables: Table 11(a), Table 11(b), and Table 11(c). These tables individually display the error rates for gibberish errors, refusal errors, and blank errors, respectively, across all tested models and languages in the 2-shot en-x xMRC task setting.
@@ B.4 Detailed language error and generation error rates
DeepSeek-V3 0.00 1.84 0.94 0.43 0.49 1.72 2.29 1.48 1.40 3.09 0.84 3.08
(a) Qwen-judged results
LLaMA-3.1-8B 4.80 4.85 7.53 7.61 4.18 6.06 5.88 7.77 7.70 6.44 5.98 8.31
LLaMA-3.1-70B 0.00 0.22 0.22 0.00 0.72 1.24 0.36 0.64 0.19 0.99 1.08 0.83
Mistral-V0.3-7B 0.49 1.35 0.81 4.86 4.15 9.38 6.43 5.33 1.35 0.00 9.43 7.38
Qwen-2.5-7B 1.08 1.11 0.37 0.84 0.39 0.19 0.24 1.18 0.35 1.00 0.44 1.59
Qwen-2.5-72B 0.00 0.00 0.00 0.00 0.45 0.46 0.00 1.19 0.48 0.49 0.90 0.82
DeepSeek-V2-Lite-16B 1.12 1.46 1.34 3.69 0.30 10.63 4.83 1.93 1.78 2.20 5.89 3.50
Gemma-2-9B 0.00 0.44 0.23 1.88 0.97 1.25 0.29 1.15 1.31 0.00 1.27 1.96
LLaMA-2-Chat-7B 5.67 5.92 12.93 20.38 16.46 34.95 33.13 37.56 7.78 6.40 58.56 7.92
LLaMA-3.1-Instruct-8B 0.85 0.72 1.39 1.02 0.66 0.83 1.14 1.19 2.59 1.34 0.68 1.56
LLaMA-3.1-Tuned-8B 1.85 0.71 0.34 0.36 1.33 0.31 0.78 1.05 0.64 0.34 0.76 1.50
LLaMA-3.1-Instruct-70B 2.43 1.22 1.09 1.82 1.02 0.76 1.56 0.80 3.47 0.85 1.74 2.14
Mistral-V0.3-Instruct-7B 0.55 0.41 0.40 0.38 0.00 0.35 0.38 0.72 0.39 0.85 0.00 0.00
Qwen-2.5-Instruct-7B 0.00 0.00 0.27 0.00 0.00 0.85 0.63 0.54 1.11 0.55 0.52 0.75
Qwen-2.5-Instruct-72B 0.58 2.33 0.84 2.87 0.93 2.43 1.89 2.41 4.01 4.21 1.49 1.52
DeepSeek-V2-Chat-Lite-16B 0.65 0.90 0.88 0.48 0.00 0.00 0.82 0.00 1.47 0.95 0.47 1.26
Gemma-2-IT-9B 0.57 0.32 0.78 0.32 0.87 1.20 1.82 0.72 0.00 0.93 0.78 1.69
GPT-3.5-Turbo-0125 0.00 0.00 0.00 0.52 0.00 0.00 0.00 0.00 0.89 0.00 0.00 0.00
GPT-4o 0.58 0.92 0.47 0.00 0.00 0.43 0.46 0.00 0.47 1.33 0.42 0.44
DeepSeek-V3 0.44 0.70 0.58 1.26 0.61 0.84 0.54 0.28 0.91 0.62 0.62 0.61
(b) Gemini-judged results
Table 10: 2-shot generation error rates (%) on en-x xMRC tasks judged by Qwen and Gemini.
@@
Table 11: Detailed percentages (%) of generation error types (gibberish error, refusal error, and blank error) on 2-shot en-x tasks.
Appendix C Two-phased xMRC Analysis
C.1 Further analysis on MRD
C.1.1 Example of attribution results
@@
C.1.2 MRD for other LLaMA models
Figures 10, 11 and 12 provide further illustrative examples of the mean MRD for context and question components, specifically for LLaMA-3.1-8B, LLaMA-3.1-Instruct-70B and LLaMA-2-Chat-7B. These figures complement the MRD analysis presented in the main body of this paper.
@@ C.1.2 MRD for other LLaMA models
C.1.3 Analysis of task descriptions and demonstrations
Analyzing the MRD of task descriptions and demonstrations in our 2-shot setting (Figures 13-15) reveals a general trend where demonstrations tend to exhibit a comparable or slightly higher MRD than task descriptions across the LLaMA model family, suggesting demonstrations are at least as important as, if not slightly more impactful than, task descriptions in guiding the models. This could indicate that providing concrete examples is a particularly effective way to communicate the desired behavior for cross-lingual context retrieval to these models.
@@ C.1.3 Analysis of task descriptions and demonstrations
C.1.4 Influence of prompt format on MRD pattern
@@
Figure 16: Mean MRD for LLaMA-3.1-Instruct-8B on both “balanced” and “en-superior” samples in v1 prompting format. Only the results of context and question parts of the prompt are displayed.
C.2 Hidden state similarity results for other LLaMA models
@@
C.3 Training details and evaluation results of our finetuned LLaMA-3.1-8B